OpenStudy (richyw):

Trying to find \(\sin^{-1}(-2)\)

OpenStudy (anonymous):

arcsine has the domain of [-1,1]

OpenStudy (loser66):

I tried very hard too, and could not get either. :)

OpenStudy (primeralph):

Nope @sourwing . Complex answers exist.

OpenStudy (richyw):

I know that I can plug it into\[\sin^{-1}z=-i\log\left[iz+(1-z^2)^{1/2}\right]\]

OpenStudy (primeralph):

So plug in.

OpenStudy (richyw):

but I am supposed to do it by writing \(\sin w = -2\) and using the rectangular representation of \(\sin w\)

OpenStudy (anonymous):

@primeralph well, if complex number is allowed then, ok. I was referring to real numbers only.

OpenStudy (primeralph):